OpenStudy (anonymous):

Rewrite the expression using a radical. \[(4+x)^{\frac{3}{2}}\]

OpenStudy (anonymous):

Remember when you have a fraction exponent like this: \[x^{\frac{5}{3}} = \sqrt[3]{x^{5}}\]?

OpenStudy (anonymous):

Yes...

OpenStudy (anonymous):

\[\sqrt{(4+x)^3}\]

OpenStudy (anonymous):

Is that it?

OpenStudy (anonymous):

Yup :)

OpenStudy (anonymous):

Do I have to expand it? Or how do I got about expanding it @Calcmathlete

OpenStudy (anonymous):

I think that's fine? You can expand if you want, but I don't think so...

OpenStudy (anonymous):

To expand it, you do this \[\sqrt{(4 + x)(4 + x)(4 + x)}\]and multiply polynomials.

OpenStudy (anonymous):

There is no way I can go straight to the final answer?

OpenStudy (anonymous):

\[\sqrt{4^3+48x+12x^2+x^3}\]

OpenStudy (anonymous):

I got \[\sqrt{x^{3} + 12x^{2} + 48x + 64}\]

OpenStudy (anonymous):

which is the same thing :).

OpenStudy (anonymous):

Thanks for all your help!

OpenStudy (anonymous):

np :)